1. Given the utility fucntion u =150x +40x2-x3derive average and marginal utility functions, Find the value of X at which total utility is maximum, and the value of X at which average utility is maximinum?
1
Expert's answer
2021-11-22T10:01:23-0500

Average utility;

Lets assume number of units consumed is "x"

Average utility=total utility"\div" no of units consumed

="\frac{150x+40x^2-x^3}{x}"

Marginal utility;

Marginal utility, MU="\frac{\Delta u}{\Delta x}=150+80x-3x^2"


Total utility;

Total utility is maximized when MU=0

"150+80x-3x^2=0"

D = 8,200

"x_{1} = \frac{-80 + 8,200^{0.5}} {2×3} = 1.76."

"x_{2}" < 0, so is not suitable for our case


Value of x at which average utility is maximum;

="\frac{150x+40x^2-x^3}{x}"

"150+40x-x^2=0"

"x=\frac{−40±√2200}{−2}"

"x_{1}=20\u22125\u221a22 =-3.45"

"x_{2}=20+5\u221a22=43.45"

Our value will be 43.45